# From Spatial to Actions: Grounding Vision-Language- Action Model in Spatial Foundation Priors

### v1

@@ 1 Introduction
Recent advances in vision-language-action models (VLAs) have significantly advanced the pursuit of generalist robotics, enabling robots to interpret natural language instructions and execute intricate action sequences [5, 13, 3, 17]. Most advanced VLAs are built on 2D foundation models like Vision Language Models (VLMs) [22, 32, 1, 30, 28, 21], aligning 2D images with text and leveraging the strong language understanding of the Large Language Models (LLMs) for information processing. This design provides strong semantic understanding and supports manipulation tasks conditioned on language and camera inputs.
Figure 1: We propose FALCON, a vision-language-action model that achieves robust 3D spatial understanding by seamless integrating spatially rich tokens and sematic features. FALCON demonstrates exceptional modality transferability by excelling with both RGB-only and multi-modal inputs, superior spatial understanding in tasks involving unseen object sizes, heights and abstract spatial instructions, and strong few-shot generalizability in real-world scenes. The model achieves state-of-the-art performance across a diverse range of benchmark evaluations.
However, while VLMs operate purely in the 2D domain, VLAs must interact with the 3D physical world. This discrepancy results in a critical gap: current VLAs lack reliable 3D spatial understanding, leading to persistent challenges in generalization and adaptability. Specifically, the absence of explicit 3D awareness causes VLAs to struggle in scenarios that require reasoning about geometry, depth, or spatial relations. First, they exhibit limited generalization, failing to transfer robustly to novel scenes, backgrounds, or object variations [42]. Second, they lack adaptability to environmental changes, such as height variations or object scale differences [16]. These limitations now form a major bottleneck to developing reliable generalist robotic policies.
@@
Contribution. We propose FALCON (From Spatial to Actions), a novel paradigm that injects richer and more representative 3D spatial tokens into VLAs through an improved injection scheme.
To solve limitation (1) i.e., weak 3D spatial representation, we leverage insights from spatial foundation models that encode scenes into token sequences for holistic 3D reconstruction. So FALCON adopts broader and richer tokens from these foundation models, and delivers comprehensive spatial information from RGB signals alone, improving robustness when explicit 3D sensors are absent.
For limitation (2) of poor modality transferability, we introduce an Embodied Spatial Model that can optionally integrate extra 3D modalities (e.g., depth, poses). Unlike prior methods that mandate specific 3D inputs [16, 33, 44], our design is flexible: when RGB-D cameras or calibrated scenes are available, the model gains additional accuracy; when not, it still performs effectively with RGB-only input. This optional pathway substantially improves transferability by natively absorbing 3D priors from sensors without requiring architectural changes.
To overcome limitation (3) of alignment challenges, we draw inspiration from the brain’s division of labor. The VLM (cerebrum) handles high-level reasoning and semantics, while the action head (cerebellum) manages fine-grained motor control and sensorimotor integration [31, 10]. Motivated by this, we design a Spatial-Enhanced Action Head that directly incorporates spatial tokens into action decisions—a more natural fit, since precise control depends on detailed spatial cues. This departs from prior approaches that forcibly align spatial and text tokens within VLMs [9, 39].
In this way, FALCON provides (i) robust spatial reasoning, (ii) strong modality transferability, and (iii) principled integration of 3D priors into VLAs. As shown in Fig. 1, extensive experiments across three simulation benchmarks and 11 real-world tasks—including cluttered-scene manipulation, few-shot adaptation, and spatial capability evaluation—demonstrate that FALCON consistently outperforms existing baselines, achieving state-of-the-art performance with strong robustness and generalization (e.g., spatial-prompt conditioning, height-aware manipulation, object-scale variation).
@@ 2 Related Work
The development of generalist robot policies has been significantly advanced by VLAs [5, 13, 18, 17, 6, 3, 14, 7, 43], which leverage large-scale pre-trained VLMs to connect visual and linguistic understanding with action generation. For instance, RT-2 [5] and OpenVLA [13] fine-tune VLMs on robot data, representing actions as language tokens. While demonstrating impressive instruction-following capabilities, these methods operate primarily in the 2D domain, lacking explicit mechanisms for 3D geometric perception, which is a critical limitation in tasks requiring precise spatial understanding. To address this, recent works have integrated 3D perceptual cues into policy learning. One line of research directly consumes explicit 3D representations like point clouds for action prediction, as seen in PointVLA [16] and GeoVLA [33]. While enhancing geometric awareness, these methods rely on specific 3D sensor configurations, limiting modality transferability when such inputs are unavailable. An alternative strategy embeds 3D features into the VLM’s input space, exemplified by 3D-VLA [44], SpatialVLA [29], and Evo-0 [20], but this often disrupts pre-trained vision-language alignment, necessitating costly embodied instruction tuning to recover performance. FALCON overcomes these issues by decoupling spatial processing from the VLM, preserving its semantic integrity while achieving strong modality transferability and robust performance across diverse input conditions.
@@ 3 Methodology
In this section, we first formulate the problem of task-oriented, language-guided robot control (Sec. 3.1). We then introduce the overall architecture of FALCON (Sec. 3.2), and detail its two key components: the Embodied Spatial Model (ESM) for 3D geometric representation (Sec. 3.3), and the Spatial-Enhanced Action Head for multimodal fusion and action generation (Sec. 3.4).
@@ 3.1 Problem Definition
$A_{t}=\mathcal{F}(O_{t},L,D_{t},P).$ (1)
@@ 3.2 Overall Architecture
We provide detailed training paradigms of the FALCON model in Appendix 6.
@@ 3.3 Embodied Spatial Model
3D Conditions Encoding. First, as shown in Fig. 2, given camera pose $P\in\mathbb{R}^{7}$ input, we encode the intrinsic and normalized extrinsic of the side camera into the GT camera token $\textbf{t}_{\text{gt-cam}}\in\mathbb{R}^{D_{\text{s}}}$ using an MLP-based camera encoder $\mathcal{E}_{cam}(\cdot)$. Given depth $D_{t}\in\mathbb{R}^{H\times W}$ and its valid map $M_{\text{dpt}}\in\mathbb{R}^{H\times W}$ input, we first normalize depth $D^{\prime}_{t}=D_{t}/\text{Norm}(D_{t})$ to handle any depth ranges at train and test time. The normalized depth map $D^{\prime}_{t}$ is concatenated with its corresponding validity mask $M_{\text{dpt}}$, enabling us to capture frame-wise incomplete depth information. The resulting tensor is then passed through a depth encoder $\mathcal{E}_{\text{dpt}}(\cdot)$, which consists of a stack of convolutional layers with a kernel size $14\times 14$, effectively partitioning it into a sequence of tokens $\mathbf{T}_{\text{dpt}}$ that are aligned in size with the image tokens $\mathbf{T}_{\text{spl}}$. The formulation is shown below:
@@ 3.4 Spatial-Enhanced Action Head
Modality Fusion Strategies. To combine these complementary representations, we first compress the spatial tokens $\mathbf{T}_{\text{spl}}$ into a unified vector $\mathbf{t}_{\text{spl}}\in\mathbb{R}^{D_{\text{s}}}$ through a max-pooling operation, then project it into the VLM’s feature space to obtain $\widetilde{\mathbf{t}}_{\text{spl}}$ using a lightweight MLP adapter $\mathcal{D}$. The aligned spatial feature $\widetilde{\mathbf{t}}_{\text{spl}}$ is then fused with the semantic action token $\hat{\mathbf{t}}_{\text{act}}$ through:
As illustrated in Fig. 3, empirical evaluation on the manipulation benchmark demonstrates that while cross-attention and FiLM-gated offer richer interaction mechanisms, element-wise addition achieves the highest performance in both seen and unseen environment generalization settings (see Sec. 4.3 for detailed quantitative results and analysis). This approach not only yields superior success rates but also provides greater training stability and computational efficiency. Thus, element-wise addition is adopted to fuse the aligned spatial feature $\widetilde{\mathbf{t}}_{\text{spl}}$ with the semantic action token $\hat{\mathbf{t}}_{\text{act}}$. The overall fusion process is formulated as:
@@ 4.2 Real-World Experiments
Base Tasks. As shown in Fig. 4, FALCON achieves the highest average success rate of 70.0% across all nine task suites, outperforming the advanced method SpatialVLA (44.4%) by 25.6%. Moreover, in the task pick banana and place on red plate, while RoboVLM and OpenVLA-OFT often erroneously place the banana onto a yellow plate, FALCON consistently places it on the correct red plate, demonstrating precise instruction following capability and superior scene understanding.
@@ 4.3 In-Depth Analysis
[…]
Table 4: Performance comparison of different modality input on CALVIN benchmark. Kosmos-VLA (w/ rgb-d) is a point cloud-based variant where the ESM is replaced by a lightweight point cloud encoder [42] while retaining other components.
@@
Figure 7: Performance comparison of different modality input on real-world tasks. Left: lift yellow pepper. Right: put white cup on pink cloth-cup height change.
Embodied Spatial Model. To further investigate the role of ESM, we conducted additional experiments on CALVIN to evaluate the monocular depth results of ESM. These experiments further validate why ESM effectively leverages additional modality inputs to achieve performance gains. As shown in Tab. 5, our ESM achieves performance comparable to VGGT when using only RGB input. Moreover, its performance improves significantly when additional depth or camera pose information is accessible. This demonstrates the inherent strength of FALCON’s modality transferability, which stems from the ability of our proposed ESM to seamlessly benefit from diverse 3D modality inputs. Fig. 8 presents ESM predicted depth maps and corresponding error maps (darker colors indicate smaller errors) under different input modalities.
We provide further ablation studies in Appendix 10 to validate key design choices in FALCON, including spatial token injection strategies and fusion mechanisms.
Table 5: Ablation study on modality inputs for ESM on the CALVIN benchmark.
Method Depth Camera $\delta<1.25$ (%) $\uparrow$ Abs. Rel $\downarrow$
VGGT [36] - - 91.33 8.53
@@
Figure 8: Depth Visualization.
@@ 5 Conclusion
In this work, we introduce FALCON, a vision-language-action model that augments generalist robot policies with robust 3D spatial understanding. FALCON makes three main contributions: (1) integration of spatial tokens from foundation models to provide strong geometric priors; (2) an Embodied Spatial Model that optionally incorporates 3D modalities (e.g., depth, camera poses) while preserving RGB-only functionality; and (3) a Spatial-Enhanced Action Head that injects spatial tokens directly into the control policy, avoiding disruptive alignment within the VLM. Experiments across both simulation and real-world tasks show that FALCON consistently surpasses existing VLA methods, achieving state-of-the-art performance and robustness on spatially demanding tasks.
@@ Acknowledgements
We would like to thank all members at ByteDance Seed Robotics team for their support throughout this project. We also want to extend our gratitude to Wenke Xia for his essential assistance with the robotic hardware setup and to Hang Li for his leadership of this team.
@@ 6.1 FALCON Training Paradigm
To integrate 3D spatial awareness into the pre-trained VLA model while preserving its generalizable action capabilities acquired during pre-training, we carefully design a two-stage post-training pipeline that effectively incorporates spatial tokens from the ESM $\mathcal{G}$. Rather than incorporating spatial information during pre-training—which would significantly increase computational cost and complicate optimization—we adopt a post-training approach that preserves the original pre-training efficiency while enabling seamless integration of 3D geometric cues.
Let $\Theta_{V}$, $\Theta_{A}$, $\Theta_{G}$, and $\Theta_{D}$ denote the parameter sets of the VLM $\mathcal{V}$, action head $\pi$, ESM $\mathcal{G}$, and lightweight adapter $\mathcal{D}$, respectively. The overall training objective is to minimize the expected action prediction loss $\mathcal{L}$ (refer to Eq. 6) over the target data distribution:
@@ 6.2 Embodied Spatial Model Training Paradigm
ESM is trained following the dataset and preprocessing procedures of VGGT [36]. For each training batch, we randomly sample 1–12 frames from a randomly selected scene, resulting in a total of 24 images per batch. Training is performed using the AdamW optimizer with differentiated learning rates: 1e-6 for the large unified transformer backbone and 1e-5 for the depth, camera, and point heads. The complete training process requires 16 A100 GPUs and runs for approximately 2 days.

### v2

@@ 1 Introduction
Recent advances in vision-language-action models (VLAs) have significantly advanced the pursuit of generalist robotics, enabling robots to interpret natural language instructions and execute intricate action sequences [5, 14, 3, 18, 48, 12, 24]. Most advanced VLAs are built on 2D foundation models like Vision Language Models (VLMs) [23, 34, 1, 32, 30, 22], aligning 2D images with text and leveraging the strong language understanding of the Large Language Models (LLMs) for information processing. This design provides strong semantic understanding and supports manipulation tasks conditioned on language and camera inputs.
Figure 1: We propose FALCON, a vision-language-action model that achieves robust 3D spatial understanding by effectively integrating spatially rich tokens and semantic features. FALCON demonstrates notable modality transferability by performing robustly with both RGB-only and multi-modal inputs, superior spatial understanding in tasks involving unseen object sizes, heights and abstract spatial instructions, and strong few-shot generalizability in real-world scenes. The model achieves state-of-the-art performance across a diverse range of benchmark evaluations.
However, while VLMs operate purely in the 2D domain, VLAs must interact with the 3D physical world. This discrepancy results in a critical gap: current VLAs lack reliable 3D spatial understanding, leading to persistent challenges in generalization and adaptability. Specifically, the absence of explicit 3D awareness causes VLAs to struggle in scenarios that require reasoning about geometry, depth, or spatial relations. First, they exhibit limited generalization, failing to transfer robustly to novel scenes, backgrounds, or object variations [46]. Second, they lack adaptability to environmental changes, such as height variations or object scale differences [17]. These limitations now form a major bottleneck in developing reliable generalist robot policies.
@@
Contribution. We propose FALCON (From Spatial to Action), a novel paradigm that integrates richer and more representative 3D spatial tokens into VLAs through an improved injection scheme.
To solve limitation (1) i.e., weak 3D spatial representation, we leverage insights from spatial foundation models that encode scenes into token sequences for holistic 3D reconstruction. FALCON adopts broader and richer tokens from these foundation models, and delivers comprehensive spatial information from RGB signals alone, improving robustness when explicit 3D inputs are absent.
For limitation (2) of poor modality transferability, we introduce an Embodied Spatial Model that can optionally integrate extra 3D modalities (e.g., depth, poses). Unlike prior methods that require specific 3D inputs [17, 37, 49], our design is flexible: when RGB-D cameras or calibrated scenes are available, the model gains additional accuracy; when not, it still performs effectively with RGB-only input. This optional pathway substantially improves transferability by flexibly utilizing 3D signals from sensors without requiring architectural changes.
To overcome limitation (3) of alignment challenges, we draw inspiration from the brain’s division of labor. The VLM (cerebrum) handles high-level reasoning and semantics, while the action head (cerebellum) manages fine-grained motor control and sensorimotor integration [33, 10]. Motivated by this, we design a Spatial-Enhanced Action Head that directly incorporates spatial tokens into action decisions, which is a more natural fit since precise control depends on detailed spatial cues. This departs from prior approaches that forcibly align spatial and text tokens within VLMs [9, 43].
In this way, FALCON provides (i) robust spatial reasoning, (ii) strong modality transferability, and (iii) principled integration of 3D priors into VLAs. As shown in Fig. 1, extensive experiments across three simulation benchmarks and 11 real-world tasks which including cluttered-scene manipulation, few-shot adaptation, and spatial capability evaluation, demonstrate that FALCON consistently outperforms existing baselines, achieving state-of-the-art performance with strong robustness and generalization (e.g., spatial-prompt conditioning, height-aware manipulation, object-scale variation).
@@ 2 Related Work
The development of generalist robot policies has been significantly advanced by VLAs [5, 14, 19, 18, 6, 3, 15, 7, 47, 35, 36], which leverage large-scale pre-trained VLMs to connect visual and linguistic understanding with action generation. For instance, RT-2 [5] and OpenVLA [14] fine-tune VLMs on robot data, representing actions as language tokens. While demonstrating impressive instruction-following capabilities, these methods operate primarily in the 2D domain, lacking explicit mechanisms for 3D geometric perception, which is a critical limitation in tasks requiring precise spatial understanding. To address this, recent works have integrated 3D perceptual cues into policy learning. One line of research directly consumes explicit 3D representations like point clouds for action prediction, as seen in PointVLA [17] and GeoVLA [37]. While enhancing geometric awareness, these methods rely on explicit 3D inputs and specific sensor configurations, limiting modality transferability when such inputs are unavailable. An alternative strategy embeds 3D features into the VLM’s input space, exemplified by 3D-VLA [49], SpatialVLA [31], and Evo-0 [21], but this often disrupts pre-trained vision-language alignment, necessitating costly embodied instruction tuning to recover performance. FALCON overcomes these issues by decoupling spatial processing from the VLM, preserving its semantic integrity while achieving strong modality transferability and robust performance across diverse sensory conditions, reducing dependency on specialized sensor setups when deploy in the real-world.
@@ 3 Methodology
In this section, we first formulate the problem of task-oriented, language-guided robot control (Sec. 3.1). We then introduce the overall architecture (Sec. 3.2), training objective (Sec. 3.3) of FALCON, and detail its two key components: the Embodied Spatial Model (ESM) for 3D geometric representation (Sec. 3.4), and the Spatial-Enhanced Action Head for multimodal fusion and action generation (Sec. 3.5).
@@ 3.1 Problem Definition
$A_{t}=\mathcal{F}(O_{t},L,D_{t},P),\quad\text{where }D_{t}\text{ and }P\text{ are optional}.$ (1)
@@ 3.2 Overall Architecture
3.3 Training Objective
To integrate spatial awareness into the VLA model while preserving the pre-trained components’ capabilities, we carefully design a two-stage post-training pipeline. Drawing inspiration from [22], in Stage 1 we freeze all pre-trained components and optimize only a lightweight adapter to achieve initial alignment between spatial tokens and the VLA’s feature space. Building upon this aligned foundation, Stage 2 unfreezes both the VLM and adapter for joint refinement while keeping the remaining parts frozen, thereby enabling the VLM implicitly refines it’s semantic features to incorporate spatial cues, which subsequently inform action prediction. The detailed training paradigms of the FALCON model are available in Appendix 6.
@@ 3.4 Embodied Spatial Model
3D Conditions Encoding. First, as shown in Fig. 2, given camera pose $P\in\mathbb{R}^{7}$ input, we encode the intrinsic and normalized extrinsic of the side camera into the GT camera token $\textbf{t}_{\text{gt-cam}}\in\mathbb{R}^{D_{\text{s}}}$ using an MLP-based camera encoder $\mathcal{E}_{cam}(\cdot)$. Given depth $D_{t}\in\mathbb{R}^{H\times W}$ and its valid map $M_{\text{dpt}}\in\mathbb{R}^{H\times W}$ input, we first normalize depth $D^{\prime}_{t}=D_{t}/\text{Norm}(D_{t})$ to handle any depth ranges at train and test time. The normalized depth map $D^{\prime}_{t}$ is concatenated with its corresponding validity mask $M_{\text{dpt}}$, enabling us to capture frame-wise incomplete depth information. The resulting tensor is then passed through a depth encoder $\mathcal{E}_{\text{dpt}}(\cdot)$, which consists of a stack of convolutional layers with a kernel size $14\times 14$, effectively partitioning it into a sequence of tokens $\mathbf{T}_{\text{dpt}}$ that are aligned in size with the image tokens $\mathbf{T}_{\text{vis}}$. The formulation is shown below:
@@ 3.5 Spatial-Enhanced Action Head
Modality Fusion Strategies. To combine these complementary representations, we first compress the spatial tokens $\mathbf{T}_{\text{spl}}$ into a unified vector $\mathbf{t}_{\text{spl}}\in\mathbb{R}^{D_{\text{s}}}$ through a max-pooling operation, then project it into the VLM’s feature space to obtain $\widetilde{\mathbf{t}}_{\text{spl}}$ using a lightweight MLP adapter $\mathcal{D}$. The aligned spatial feature $\widetilde{\mathbf{t}}_{\text{spl}}$ is then fused with the semantic action token $\hat{\mathbf{t}}_{\text{act}}$ through highly efficient element-wise addition (see Sec. 4.3 for detailed ablations about different fusion strategies). The overall fusion process is formulated as:
@@ 4.2 Real-World Experiments
Base Tasks. As shown in Fig. 3, FALCON achieves the highest average success rate of 70.0% across all nine task suites, outperforming the advanced method SpatialVLA [31] (44.4%) by 25.6%. Moreover, in the task pick banana and place on red plate, while RoboVLM [18] and OpenVLA-OFT [15] often erroneously place the banana onto a yellow plate, FALCON consistently places it on the correct red plate, demonstrating precise instruction following capability and superior scene understanding.
@@ 4.3 In-Depth Analysis
Table 4: Ablation studies on spatial token injection methods and fusion strategies. Results confirm that the standard FALCON paradigm achieves the best performance, validating it as the optimal design choice.
@@
[…]
Modality Fusion Strategies. To combine the aligned spatial feature $\widetilde{\mathbf{t}}_{\text{spl}}$ with the semantic action token $\hat{\mathbf{t}}_{\text{act}}$, we further investigate three distinct fusion approaches:
As shown in Tab. 4, element-wise addition (our standard FALCON) consistently delivers the best performance across all experimental settings. This method achieves the highest task success rates while remaining both simple and computationally efficient, as it introduces no additional parameters. The results underscore the effectiveness of a straightforward, parameter-free fusion strategy for combining spatial and semantic representations in VLA models.
Table 5: Performance comparison of different modality input on CALVIN benchmark. Kosmos-VLA (w/ rgb) is a 2D VLA without ESM. Kosmos-VLA (w/ rgb-d) is a point cloud-based variant where the ESM is replaced by a lightweight point cloud encoder [46] while retaining other parts.
Method Task Tasks Completed in a Row (%) Avg. Len. $\uparrow$
1 2 3 4 5
@@
[…]
Figure 7: Performance comparison of different modality input on real-world tasks. Left: lift the yellow pepper. Right: put white cup on pink cloth–cup height change.
[…]
Embodied Spatial Model. To further investigate the role of ESM, we conduct additional zero-shot experiments on CALVIN to evaluate the monocular depth estimation results of ESM. These experiments further validate why ESM effectively leverages additional modality inputs to achieve performance gains. As shown in Tab. 6, our ESM achieves performance comparable to VGGT when using only RGB input. Moreover, its performance improves significantly when additional depth or camera pose information is accessible. This demonstrates the inherent strength of FALCON’s modality transferability, which stems from the ability of our proposed ESM to seamlessly benefit from diverse 3D modality inputs. Fig. 8 presents ESM predicted depth maps and corresponding error maps (darker colors indicate smaller errors) under different input modalities.
Table 6: Zero-shot monocular depth estimation with diverse input modalities for ESM on CALVIN benchmark. $\delta<1.25$: measures the accuracy under 1.25 threshold (percentage of predictions with relative error $\leq 25\%$). Abs. Rel: computes the average absolute relative error between prediction and ground truth.
Method Depth Camera $\delta<1.25$ (%) $\uparrow$ Abs. Rel $\downarrow$
VGGT [40] - - 91.33 8.53
Ours 90.91 8.61
Ours 99.79 0.91
Ours 99.47 0.87
Figure 8: Depth map Visualization.
@@ 5 Conclusion
In this work, we introduce FALCON, a vision-language-action model that augments generalist robot policies with robust 3D spatial understanding. FALCON makes three main contributions: (1) integration of spatial tokens from foundation models to provide strong geometric priors; (2) an Embodied Spatial Model that optionally incorporates 3D modalities (e.g., depth, camera poses) while preserving RGB-only functionality; and (3) a Spatial-Enhanced Action Head that injects spatial tokens directly into the control policy, avoiding disrupting vision-language alignment within the VLM. Experiments across both simulation and real-world tasks show that FALCON consistently surpasses existing VLA methods, achieving state-of-the-art performance and robustness on spatially demanding tasks.
@@ Acknowledgements
We would like to thank all members at ByteDance Seed Robotics team for their continuous support throughout this project. We also want to extend our gratitude to Wenke Xia for his essential assistance with the robotic hardware setup, to Haosong Peng for his valuable discussions, and to Hang Li for his leadership of this team.
Ethics Statement
This work focuses on improving spatial reasoning for robotic manipulation. While our model is trained on standard open-source datasets and tested in controlled settings, we acknowledge that any AI system may potentially exhibit biases or produce unexpected behaviors. Our research is intended for academic exploration only, and we emphasize that any such outcomes do not reflect the views of the authors. We support the development of AI technologies that are ethical, safe, and aligned with societal values.
Reproducibility Statement
We detail our work in the Methodology section (Sec. 3) and describe implementation details at the beginning of Sec. 4 and Appendix 6, 7, 8, 9.
Declaration of LLM Usage
In the preparation of this work, the authors used LLM (e.g., GPT-4) in order to improve the readability and language of the manuscript. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
@@ 6.1 FALCON Training Paradigm
Rather than incorporating spatial information during pre-training, which would significantly increase computational cost and complicate optimization, we adopt a post-training approach that preserves the original pre-training efficiency while enabling efficient integration of 3D geometric cues.
Let $\Theta_{V}$, $\Theta_{A}$, $\Theta_{G}$, and $\Theta_{D}$ denote the parameter sets of the VLM $\mathcal{V}$, action predictor $\pi$, ESM $\mathcal{G}$, and lightweight adapter $\mathcal{D}$, respectively. The overall training objective is to minimize the expected action prediction loss $\mathcal{L}$ (refer to Eq. 2) over the target data distribution:
@@ 6.2 Embodied Spatial Model Training Paradigm
ESM is trained following the dataset and preprocessing procedures of VGGT [40]. For each training batch, we randomly sample 1–12 frames from a randomly selected scene, resulting in a total of 24 images per batch. Training is performed using the AdamW optimizer with differentiated learning rates: 1e-6 for the large unified transformer backbone and 1e-5 for the depth, camera, and point heads. We set Bernoulli probability $p=66\%$ throughout ESM training. The complete training process requires 16 A100 GPUs and runs for approximately 2 days.